Arvidas recently won the Central States Lottery of $3,500,000. The lottery pays either a total of twenty $175,000 payments per year with the first payment today (i.e., an annuity due), or $2,000,000 today. At what interest rate would Arvidas be financially indifferent between these two payout choices?

8.00%

6.88%

6.04%

6.36%
Answer #1

Ans 6.88%

Annuity PV Due (Beginning)= P + P [ 1 - ( 1 + r )-(n-1) ]
                       r
2000000 = 175000 + 175000 * ( 1 - ((1 / (1 + r%)^(20-1))))/ (r%)
(2000000 - 175000 ) / 175000 = ( 1 - ((1 / (1 + r%)^(20-1))))/ (r%)
10.4285714285714 = ( 1 - ((1 / (1 + r%)^(20-1))))/ (r%)
r = 6.88%